OpenStudy (anonymous):

Identify the roots of these equation. 1)x^3-5x^2=8x-4=0 2)x^4-12x^2+27=0

OpenStudy (anonymous):

graph them

OpenStudy (anonymous):

isn't there another way to do it too?

OpenStudy (anonymous):

not that i know of

OpenStudy (anonymous):

I don't have a graphing calculator at school though.. dang it. thanks though

OpenStudy (anonymous):

u dont n ur welcome

OpenStudy (anonymous):

retype number 1) please since there's an equal sign in there.

OpenStudy (anonymous):

Oh sorry @OrionsBelt 1)x^3-5x^2+8x-4=0

OpenStudy (anonymous):

i got 1 & 2

OpenStudy (anonymous):

I have to show work, can someone please explain how to do these probelsm with work? Please?

OpenStudy (anonymous):

if i knew how to do it with work i would tell u but i graphed it on the calculator

OpenStudy (anonymous):

it's fine- My teacher wants me to show work though that's the problem haha

OpenStudy (anonymous):

Your second equation. Use u = x^2 as the substitution. Then it should give you a quadratic equation. From there solve the equation then plug back in your u = x^2 then solve for "x"

OpenStudy (anonymous):

wait what do you mean by use u=x^2 as the substitution?

OpenStudy (anonymous):

This is your 2nd equation: \[x^4-12x^2+27 = 0\] and we use the letter "u" and set it equal to "x^2" then sub it into the equation so it turns out as so: \[u^2 -12u +27 =0\]You can solve the equation from here for the quadratic equation

OpenStudy (anonymous):

for your first equation look at this link: http://answers.yahoo.com/question/index?qid=20120304065333AA4OpJL

OpenStudy (anonymous):

ok,so- for the second equation, for the quadratic Formula- a=1,b=12 and c=27.. correct?

OpenStudy (anonymous):

No. Put that quadratic equation into factored form, do you know how to do that?

OpenStudy (anonymous):

I might,but I don't remember exactly

OpenStudy (anonymous):

Once you put your "u" sub in you get this original equation and have to solve for quadratic equation that equals zero as so: \[u^2-12u+27 =0\]\[(u-9)(u-3)=0\]your equation is still in "u" form. We want it in "x" form which we plug back in our original "x" which we had u=x^2 right? so the final equation will look like this: \[(x^2-9)(x^2-3)=0\] then you solve each them separately by setting them both equal to zero as so:\[x^2-9=0\]\[x^2-3 =0\]you can solve the rest from here right?

OpenStudy (anonymous):

yeah add 9 and 3 to both sides and then square both sides, right?

OpenStudy (anonymous):

yes square root both sides.

OpenStudy (anonymous):

alright thankyou!